Newton Industries is considering a project and has developed the following estimates: unit sales = 7,300, price per unit = $149, variable cost per unit = $91, fixed costs = $216,400. The depreciation is $94,700 a year and the tax rate is 35 percent. What effect would an increase of $1 in the selling price have on the operating cash flow?
1
Expert's answer
2012-11-06T09:33:51-0500
Initial operating Cash Flow is
(7300*149-91*7300-216400)*(1-0,35)+94700= (1087700-664300-216400)*0,65+94700=
229250

After price increase:
(7300*150,5-91*7300-216400)*(1-0,35)+94700= (1098577-664300-216400)*0,65+94700=
236320,05
Thus cash flow increases on 3%